# thanks: Contributed equallythanks: Contributed equally

### v2

@@ III.2 Effect of long-ranged translational repulsion
We now switch on $\chi_{t}$ and ask how the properties of the flock differ. On the (ł$(\l_{r}/b,\l_{t}/b)$ plane, we find that there is a clear region of flocking for non-zero $\chi_{t}$ (see Figure 3(a)). We further find that the flocks have crystalline order (see Figure 3(c),(e)). Such crystalline flocks in repulsive systems have been recently reported [26]. Similarly to the CLF, the CCF destabilized by sufficiently high rotational torques. In addition, $l_{+}$ now is linear in $l_{t}$. $l_{+}$ being larger for a positive $l_{t}$, this thus renders a shorter sliding length for pair colloids during the CCF formation. The net result is thus a shorter sliding length compared to the CLF, which we can attribute to the difference between crystalline and liquid structures. We note that if this repulsion is too strong, the CCF is destabilized to a disordered gas (not shown here). The most generic form of the transition lines from the phase diagrams can be written as
@@
which is presented in Figures (1)(a) and (3)(a). We obtain $\beta_{1}$ and $c_{1}$ by first fitting the boundary line in Fig. 3(a). The slope gives $\beta_{1}\approx 0.167$, whilst the $x$-intercept gives $c_{1}\approx 0.104$. This $c_{1}$ then independently fixes the boundary on Fig. 1(a). We show below how this can be derived from a continuum model of interacting density and polarization fields. The $\frac{c_{1}}{\rho}$ term promotes sliding at high densities. This effect can be understood by comparing $l_{sl}$ with the mean free path $l_{\rho}=\frac{1}{\sqrt{\rho}}$. At density values higher than $\rho_{c}=(bc_{1})^{2/3}\approx 0.058$ we have that $l_{sl}$ exceeds $l_{\rho}$. Thus, for $\rho>\rho_{c}$, sliding is exaggerated via a confinement of pair colloids due to frequent collisions. If excessive, this effect can bring $l_{sl}$ close to the upper bound $l_{sl}^{-}$, inhibiting flock formation at high densities. Instead for $\rho<\rho_{c}$, this effect is suppressed, and the pair sliding mechanism happens uninhibited, encouraging flock formation. Both these effects are consistent with our results in Fig. 1(a). This thus restricts the density range within which fluid flocks can form. For $\chi_{t}>0$, the second term further suppresses the sliding length due to long-ranged translational repulsion. Sliding over shorter distances effectively localizes particles in space and the effect is a crystalline structure.
We note that the flock destabilization mechanism here is independent of the contribution of noise. Though we have explicit noise terms in the updates in Eq.(1) for the sake of completion, they are set to be negligibly small (see Table 1 in Appendix). Indeed, if we were to define length scales $l_{\eta}=\frac{D_{r}b^{2}}{v_{s}}$ and $l_{\xi}=\frac{D_{t}}{v_{s}}$, using the values used of $D_{r}\approx 10^{-{6}}$ and $D_{t}\approx 10^{-{8}}$, we find that $l_{\eta}\sim 10^{-4}b$ and $l_{\xi}\sim 10^{-6}b$, thus negligible compared to all other length scales in the problem. We have repeated all the simulations presented in Fig. (1) with $D_{r}=D_{t}=0$, and have found the results to be identical. Thus, as opposed to the conventional picture of noise induced inhibition of flock formation [4, 11, 21, 26] (i.e. deterministic alignment interactions compete with randomness in the same interaction channel), our mechanism reports a competition between a deterministic length scale due to sensing of chemical gradients $l_{r}$ (also $l_{t}$) and other deterministic length scales of the problem.
@@ III.3 Difference between CCF and CLF
An additional feature of the CLF is that they have a steady-state persistent fluctuations of both density and polarization fields (dynamical steady-state - note that our simulations are effectively noiseless), which is absent in the CCF (see appendix B and Fig. 8). Further, we note that both the CCF and CLF show an enhanced number fluctuations $\langle(\Delta N)^{2}\rangle\sim N^{\alpha}$ with $\alpha\sim 1.2$ [1, 39, 40].
To further support the robustness of the mechanism we report CLFs arising repulsive torques with the addition of long-ranged attractive interactions, thus $\chi_{t}<0$. We find that this phenomenology is maintained for $l_{t}\in[0,-0.01]$, before transitioning to a disordered gas for larger attractive translational interactions; this is shown in Figure 3(b). Putatively, the clumping tendency in this range is balanced by the excluded volume repulsions.
Figure 4: Phase diagrams in terms of (left) global hexatic order ($\psi_{6}$) and (right) density variance. The panel on top shows results for the case (a), which is $\chi_{t}=0$, whereas the bottom panel shows results for the case (b), which is $\chi_{t}>0$. Dotted curves indicate flocking order-disorder transition line. See text for details.
@@ III.4 Hexatic order and Density variance
One defines a measure of local hexatic order $\psi_{i}$ for the $i$th particle to quantify the hexatic order globally in terms of $\psi_{6}$. They are given as [41]:
@@ III.5 CCFs as solutions to equation of motion
The hexagonal structure of CCF makes (7) tractable. Firstly, let us consider a hexagonal lattice topology of distance $2b^{\prime}$ aligned along the direction of the flock $\theta_{i}$ for the $i$th reference particle. To compute the sums in (7) one needs to evaluate the sums
@@
where $K$ is the upper bound on the radial sums performed on the lattice; in our case $K=\frac{50}{b^{\prime}}-1$. To evaluate $\prod_{k\neq i,j}|r_{i}-r_{k}|^{3}$. We can evaluate this to be $\prod_{k}^{K}((1+k)b^{\prime})^{z-1}\delta_{k,j\in\{k\}}((1+k)b^{\prime})^{z}(1-\delta_{k,j\in\{k\}})$, thus raised to power coordination number $z$ minus one for the radial coordinate containing the $j$th particle (denoted by $j\in\{k\}$), and to the coordination number for all other radial coordinates. We recognize immediately that the symmetry of the lattice structure means that this depends on only the location of the “left out” particle $j$ and the reference particle $i$. Without loss of generality, we label this product $P_{ij}$. The sums over $n$ for each component in (8) can be readily computed by applying the double angle formulae and symmetries of the sine and cosine. The overall result is thus
@@
which we see to be the self-consistent $\theta_{i}^{*}=\theta_{i}$ as required. Further, taking ratio of the coordinates of the positional dynamics at the fixed point, we see that $\frac{\dot{y}_{i}}{\dot{x}_{i}}=\tan(\theta_{i})$ as required. We thus conclude that our particular choice of chemical interactions of (3) supports a crystalline flock. For fluid flocks the result is less trivial since various sums and products need to be weighted by a distribution function. We nevertheless see that enhanced fluctuations of $\theta_{i}^{*}$ is to be expected, since any approximate evaluation of (7) will contain various moments of each components of (8).
@@ IV Continuum description
We write down here the coarse-grained hydrodynamic equations for the density $\rho(\mathbf{r},t)$ and polarization $\mathbf{p}(\mathbf{r},t)$ fields. The procedure is well established [38], and we can quote directly the equations:
@@
$v_{\text{tr}}$ $=v_{s}-\chi_{t}\xi_{0}\rho,\quad D_{\text{tr}}=\xi_{t}\rho\chi_{t},\quad\Lambda=\xi_{r}\chi_{r}\rho-v_{\text{tr}}\,\rho$
where the constants $\xi_{0}$, $\xi_{t}$, and $\xi_{r}$ are obtained by fitting the continuum model to the transition line predicted by the particle-based simulation in Fig.1. From here on we set $D_{r}$, $D_{t}$ terms to zero in (IV) as they are negligibly small as explained in the previous section. In effect, the presence of these terms merely necessitate the formal coarse-graining procedure, but do not play any further role in explaining the instability of the flock.
We wish to study the flocking state, thus we linearize (IV) about the steady-state density and polarizations ($\rho^{*},\mathbf{p}^{*}$), defining $(\delta\rho,\delta\mathbf{p})=(\rho-\rho^{*},\mathbf{p}-\mathbf{p}^{*})$. We use the following choice of Fourier basis expansion
@@ IV.1 Case of $\chi_{t}=0$
Here, we can solve equations (IV) in real space; we obtain an enslaved dynamics of $\mathbf{\dot{p}}$ (instead of the conventional $\mathbf{p}$) with respect to the density gradients:
$\partial_{t}\mathbf{p}=\frac{\nabla\rho}{2}\Big{[}(\chi_{r}\xi_{r}\rho-v_{s})\Big{]}$ (13)
@@
$\chi_{r}<\frac{v_{s}}{\xi_{r}\rho^{*}}$ (14)
We see evidence for this enslavement in Fig. (1)(a). We see that the fluctuations in both $\rho$ and $\mathbf{p}$ are strongly coupled and that any fluctuations are stabilized in the long-time limit.
@@ IV.2 General case with $\chi_{t}>0$
Firstly, we may note that solving the $\dot{\mathbf{p}}=0$ in (IV) gives $\rho^{*}=\frac{v_{s}}{\xi_{r}\chi_{r}+2\xi_{t}\chi_{t}}$, a generalization of the above inequality. Further, we need to now linearize (IV) in the bases of (12) to obtain the stability condition for the flocking state. The full linearization in Fourier space yields the following equation
@@
Here $B[\rho^{*}]=(2\xi_{0}\chi_{t}+\xi_{r}\chi_{r})\rho^{*}-v_{s}$. The dispersion relation for $\omega(q)$ then reads
@@
$\chi_{t}^{*}=\frac{v_{s}-\frac{\xi_{r}\chi_{r}\rho^{*}}{2}}{2\xi_{0}{\rho^{*}}}\Big{[}1\pm\sqrt{1-\frac{2(\frac{\xi_{r}\chi_{r}\rho^{*}}{v_{s}}-1)}{(1-\frac{\xi_{r}\chi_{r}\rho^{*}}{2v_{s}})^{2}}}\Big{]}$ (17)
Let us consider the negative solution, thus $\chi_{t}<\chi_{t}^{*}$. Let us further study the limit $\frac{\chi_{r}}{v_{s}}<<1$ (this is justified, since in simulations this value is at most $\approx 0.2$ see Table in Appendix. Using the definitions for $l_{r}$ and $l_{t}$, the desired inequality is thus
@@
We can identify this to have the same functional form as Eq.(5) defining the transition line. Equating with Eq. 4 in the main text, we have that
@@
We can thus infer the values of $\xi_{r}\approx 2c_{1}\approx 0.028$ and $\xi_{0}\approx 0.122$. We thus conclude that the noiseless limit of a minimal hydrodynamic model captures the destabilization transition of the flock. In addition, such transitions within between crystalline/disorder phases (e.g. Fig. 3(b)) are also captured, although there is no explicit spatial-structure instability encoded in the hydrodynamic model (e.g. for instance in [37]).
Figure 5: (a) Schematic of the agent-based-sliding model. (b) phase diagram in the plane of $(\rho,{l_{r}})$ from the simulations of the agent-based model (to be compared with Fig. 1(a)). Symbols represent steady-state polarization (as opposed to spatial structure) as indicated in the legend. Parameters used are $v_{s}=5$, $c_{1}=0.14$, $r_{int}=2$, $\frac{1}{\xi}=30$, with $l_{t}=0$.
V Agent-based-sliding model
To further substantiate our findings, we now construct an agent-based model of pair colloids that interact via a sliding-whilst-rotating mechanism. To make progress, we consider the dynamics (1) in the limit of $e_{x_{i}}=e_{x_{j}}$ and $e_{y_{i}}=-e_{y_{j}}$. This corresponds to any arbitrary incoming collision in the frame of $\mathbf{\hat{e}}_{sl}$, as depicted in Fig. 5(a) (the angle $\theta$ now is strictly speaking deviations from the half angle $\bar{\theta}=\frac{1}{2}(\theta_{i}+\theta_{j})$). In the sliding picture, the 4 degrees of freedom in (1) reduce to 2, for the center of mass motion $l_{sl}$ and the $i$th angle $\theta_{i}$ (we could have equivalently considered the relative angle $\Delta\theta=\theta_{i}-\theta_{j}$, and then reset $\theta_{i}=\frac{\Delta\theta}{2}$). The evolution of the latter is given by $\dot{\theta}_{i}=v_{s}\cos\theta_{i}/4l_{sl}$, with its solution given by $\frac{\theta_{i}}{2}=\tan^{-1}\Big{(}A\exp\frac{tv_{s}}{4l_{sl}}\Big{)}-\frac{\pi}{4}$, with $A=\tan(\frac{\theta_{i}^{0}}{2}+\frac{\pi}{4})$ (superscript 00 denoting the incident angle). This solution we see to be idealized since it asymptotes to $\frac{\pi}{2}$ in the long time limit. We nevertheless see that the outgoing angle grows inversely with $l_{sl}$.
To run simulations in our agent-based model, we denote $\theta^{0}_{i}$ ($\theta^{f}_{i}$) as the incident (final) angle of agent $i$ upon collision. Upon interacting with agent $j$ within an interaction radius $r_{int}$, the update equations for the model are such that whilst $||\mathbf{r}_{i}-\mathbf{r}_{j}||<r_{int}$
$\mathbf{r}_{i}^{t+1}$ $\rightarrow\mathbf{r}_{i}^{t}+l_{sl}\mathbf{{e}_{sl}},\quad\mathbf{{e}_{sl}}=\frac{1}{2}\Big{(}\mathbf{{e}}_{i}+\mathbf{{e}}_{j}\Big{)}$ (20a)
$\theta_{i}^{t+1}$ $\rightarrow\left(1-\frac{p_{sl}}{2}\right)\theta_{j}^{t}+\frac{p_{sl}}{2}\,\theta_{i}^{t}$ (20b)
with $j\leftrightarrow i$ for the $j$th agent update ($i$ and $j$ are updated simultaneously, see schematic in Fig. 5(a)). We assign the probability $p_{sl}$ of a particular slide of length $l_{sl}$ as:
$p_{sl}(l_{sl})=1-\text{erf}\left[\frac{l_{sl}}{\xi}\left(1-\frac{l_{sl}^{*}}{l_{sl}}\right)\right],\qquad l_{sl}^{*}=\frac{r_{int}^{2}}{4l_{+}}$ (21)
We note that only pair collisions are considered, whilst outside $r_{int}$, the dynamics are simply $\dot{\mathbf{r}}=v_{s}\mathbf{e}$ (noiseless self-propelled particles). As it stands, (20) does not resolve individual time steps during the collision, but consists of discrete updates during a collision event. $p_{sl}$ is constructed such that sufficiently small slides ($l_{sl}<<\frac{r_{int}^{2}}{4l_{+}}$) are more likely hence $p_{sl}\rightarrow 1$ in the limit of $\xi<<1$ (fully elastic collision, e.g. at high densities). If instead $l_{sl}>>\frac{r_{int}^{2}}{4l_{+}}$ we have an aligning collision with $\theta_{i}^{f}=\frac{1}{2}(\theta_{j}^{0}+\theta_{i}^{0})$. We note that the other extreme of $l_{r}<l_{-}$ is not considered here; since our aim is to reproduce main features of our chemical interaction-based numerical simulations. We note that the model has thus only one free parameter, $\xi$, in addition to the previously defined ones. The results of the simulation are shown in Fig. 5(b), where we see that the phase diagram of Fig. 1 is roughly reproduced. The key difference viz-a-viz Fig.1 is that in most cases a net polarization is produced, with the delimiting lines of Eq. (5) giving solutions that are fully polarized.
VI Conclusions

### v3

@@ III.2 Effect of long-ranged translational repulsion
We now switch on $\chi_{t}$ and ask how the properties of the flock differ. On the $(l_{r}/b,l_{t}/b)$ plane, we find that there is a clear region of flocking for non-zero $\chi_{t}$ (see Figure 3(a)). We further find that the flocks have crystalline order (see Figure 3(c),(e)). Such crystalline flocks in repulsive systems have been recently reported [26]. Similarly to the CLF, the CCF destabilized by sufficiently high rotational torques. In addition, $l_{+}$ now is linear in $_{t}. l+l_{+}$ being larger for a positive $l_{t}$, this thus renders a shorter sliding length for pair colloids during the CCF formation. The net result is thus a shorter sliding length compared to the CLF, which we can attribute to the difference between crystalline and liquid structures. We note that if this repulsion is too strong, the CCF is destabilized to a disordered gas (not shown here). The most generic form of the transition lines from the phase diagrams can be written as
@@
which is presented in Figures (1)(a) and (3)(a). We obtain $\beta_{1}$ and $c_{1}$ by first fitting the boundary line in Fig. 3(a). The slope gives $\beta_{1}\approx 0.167$, whilst the $x$-intercept gives $c_{1}\approx 0.013$. This $c_{1}$ then independently fixes the boundary on Fig. 1(a). We show in section IV how this can be derived from a continuum model of interacting density and polarization fields. The $\frac{c_{1}}{\rho}$ term inhibits sliding at high densities. We note that for densities above a critical density $\rho_{c}=(\frac{c_{1}}{b})^{2/3}\approx 0.05$, $l_{sl}$ always exceeds the mean free path $\frac{1}{\sqrt{\rho}}$ so that flock formation can happen uninhibited. Thus, the destabilization at higher densities occurs as $l_{sl}^{-}$ is approached due to the dominance of the instantaneous chemical sensing relative to the excluded volume repulsions (though in this regime collisions are more frequent). For $\chi_{t}>0$, the second term further suppresses the sliding length due to long-ranged translational repulsion. Sliding over shorter distances effectively localizes particles in space and the effect is a crystalline structure.
We note that the flock destabilization mechanism here is independent of the contribution of noise. Though we have explicit noise terms in the updates in Eq.(1) for the sake of completion, they are set to be negligibly small (see Table 1 in Appendix). Indeed, if we were to define length scales $l_{\eta}=\frac{D_{r}b^{2}}{v_{0}}$ and $l_{\xi}=\frac{D_{t}}{v_{0}}$, using the values used of $D_{r}\approx 10^{-{6}}$ and $D_{t}\approx 10^{-{8}}$, we find that $l_{\eta}\sim 10^{-4}b$ and $l_{\xi}\sim 10^{-6}b$, thus negligible compared to all other length scales in the problem. We have repeated all the simulations presented in Fig. (1) with $D_{r}=D_{t}=0$, and have found the results to be identical. Thus, as opposed to the conventional picture of noise induced inhibition of flock formation [4, 10, 21, 26] (i.e. deterministic alignment interactions compete with randomness in the same interaction channel), our mechanism reports a competition between a deterministic length scale due to sensing of chemical gradients $l_{r}$ (also $l_{t}$) and other deterministic length scales of the problem.
@@ III.3 Difference between CCF and CLF
To further support the robustness of the mechanism, we report CLFs arising from repulsive torques with the addition of long-ranged attractive interactions, thus $\chi_{t}<0$. We find that this phenomenology is maintained for $l_{t}\in[0,-0.01]$, before transitioning to a disordered gas for larger attractive translational interactions; this is shown in Figure 3(b). Putatively, the clumping tendency in this range is balanced by the excluded volume repulsions.
Figure 4: Phase diagrams in terms of (left) global hexatic order ($\psi_{6}$) and (right) density variance. The panel on flocktop shows results for the case (a), which is $\chi_{t}=0$ (CLF), whereas the bottom panel shows results for the case (b), which is $\chi_{t}>0$ (CCF). Dotted curves indicate flocking order-disorder transition line. See text for details.
@@ III.4 Hexatic order and density variance
We define a measure of local hexatic order $\psi_{i}$ for the $i$th particle to quantify the hexatic order globally in terms of $\psi_{6}$. They are given as [39]:
@@ III.5 CCFs as solutions to equation of motion
The hexagonal lattice topology of CCF makes (10) tractable. Firstly, let us consider a hexagonal lattice topology of distance $2b^{\prime}$ aligned along the direction of the flock $\theta_{i}$ for the $i$th reference particle. To compute the sums in (10) one needs to evaluate the sums
@@
where $K$ is the upper bound on the radial sums performed on the lattice; in our case $K=\frac{50}{b^{\prime}}-1$. For $\prod_{k\neq i,j}|r_{i}-r_{k}|^{3}$, we can evaluate this to be $\prod_{k}^{K}((1+k)b^{\prime})^{z-1}\delta_{k,j\in\{k\}}((1+k)b^{\prime})^{z}(1-\delta_{k,j\in\{k\}})$, thus raised to power coordination number $z$ minus one for the radial coordinate containing the $j$th particle (denoted by $j\in\{k\}$), and to the coordination number for all other radial coordinates. We recognize immediately that the symmetry of the lattice structure means that this depends on only the location of the “left out” particle $j$ and the reference particle $i$. Without loss of generality, we label this product $P_{ij}$. The sums over $n$ for each component in (11) can be readily computed by applying the double angle formulae and symmetries of the sine and cosine. The overall result is thus
@@
which we see to be the self-consistent $\theta_{i}^{*}=\theta_{i}$ as required. Further, taking ratio of the coordinates of the positional dynamics at the fixed point, we see that $\frac{\dot{y}_{i}}{\dot{x}_{i}}=\tan(\theta_{i})$ as required. We thus conclude that our particular choice of chemical interactions of (4) supports a crystalline flock. For fluid flocks the result is less trivial since various sums and products need to be weighted by an appropriate distribution function. We nevertheless see that enhanced fluctuations of $\theta_{i}^{*}$ is to be expected, since any approximate evaluation of (10) will contain various moments of each components of (11).
III.6 Order-disorder phase transition
Figure 5: Magnitude of the polar order parameter $|\mathbf{P}|$ versus $l_{r}/b$ in panel (a) for the case of CLF. Here $l_{t}=0$. In panel (b) we show, for the case of CCF, the variation of the order parameters as a function of $l_{t}/b$. Here, we have fixed $l_{r}=0.55$. Different symbols indicates different area fraction $\phi=\rho\pi b^{2}$.
We now present how polar order parameter $P=\langle|\mathbf{P}|\rangle_{ss}$ changes with variation of a control parameter $l_{r}/b$ and $l_{t}/b$, shown in Fig. 5. First, for CLF ($l_{t}=0$), phase transitions are shown for different densities of active particles, with area fraction $\phi=\rho\pi b^{2}$ (Fig. 5(a)). With zero $l_{r}$, $P\approx 0$ and there is no flocking (lower bound of Fig. 1(a)). As we increase $l_{r}$, particles flock into a polar ordered state. Whereas with further increase $l_{r}$, disordered phase are observed (see Fig. 1(a)). Transition points shift towards lower $l_{r}/b$ values with increase of density. Then, for the CCF (with fixed $l_{r}=0.55$), polar order parameter $P$ versus $l_{t}/b$ is plotted for different area fractions (for reference, see Fig. 4(b), where the area fraction is fixed at $\phi=0.36$). In our model, the phase transition are highly dependent on the system size and many other parameters such as $\rho$, $v_{0}$, $\chi_{t}$ and $\chi_{r}$. Though of an first-order nature, an extensive study of the exact nature of the transition is left for future work. It needs to be noted that our model includes long-range interactions, and it is computationally expensive to work with such a system.
@@ IV Continuum description
In this section, we write the coarse-grained hydrodynamic equations for the density $\rho(\mathbf{r},t)$ and polarization $\mathbf{p}(\mathbf{r},t)$ fields by appealing to the symmetries and conservation laws of the system [39]. It is also possible to obtain these equations from well-established coarse-graining procedures [17, 38]. The continuum equations for our system can be written as:
@@
$v_{\text{tr}}$ $=v_{0}-\chi_{t}\xi_{0}\rho,\quad D_{\text{tr}}=\xi_{t}\rho\chi_{t},\quad\Lambda=\xi_{r}\chi_{r}\rho-v_{\text{tr}}\,\rho$
where the constants $\xi_{0}$, $\xi_{t}$, and $\xi_{r}$ are obtained by fitting the continuum model to the transition line predicted by the particle-based simulation in Fig.1(a) and 3(a). From here on we set $D_{r}$, $D_{t}$ terms to zero in (14) as they are negligibly small as explained in the previous section. In effect, the presence of these terms merely necessitate the formal coarse-graining procedure, but do not play any further role in explaining the instability of the flock.
We now proceed to study the stability of the flocking state. To this end, we linearize (14) about the steady-state density and polarizations ($\rho^{*},\mathbf{p}^{*}$), defining $(\delta\rho,\delta\mathbf{p})=(\rho-\rho^{*},\mathbf{p}-\mathbf{p}^{*})$. We use the following choice of Fourier basis expansion
@@ IV.1 Case of $\chi_{t}=0$
Here, we can solve equations (14) in real space; we obtain an enslaved dynamics of $\mathbf{\dot{p}}$ with respect to the density gradients:
$\partial_{t}\mathbf{p}=\frac{\nabla\rho}{2}\Big{[}(\chi_{r}\xi_{r}\rho-v_{0})\Big{]}.$ (16)
@@
$\chi_{r}<\frac{v_{0}}{\xi_{r}\rho^{*}}.$ (17)
We see that the fluctuations in both $\rho$ and $\mathbf{p}$ are strongly coupled and that any fluctuations are stabilized in the long-time limit (see Appendix Fig. (7)).
@@ IV.2 General case with $\chi_{t}>0$
Firstly, we may note that solving the $\dot{\mathbf{p}}=0$ in (14) gives $\rho^{*}=\frac{v_{0}}{\xi_{r}\chi_{r}+2\xi_{t}\chi_{t}}$, a generalization of the above inequality. Further, we need to now linearize (14) in the bases of (15) to obtain the stability condition for the flocking state. The full linearization in Fourier space yields the following equation
@@
Here $B[\rho^{*}]=(2\xi_{0}\chi_{t}+\xi_{r}\chi_{r})\rho^{*}-v_{0}$. The dispersion relation for $\omega(q)$ then reads
@@
$\chi_{t}^{*}=\frac{v_{0}-\frac{\xi_{r}\chi_{r}\rho^{*}}{2}}{2\xi_{0}{\rho^{*}}}\Big{[}1\pm\sqrt{1-\frac{2(\frac{\xi_{r}\chi_{r}\rho^{*}}{v_{0}}-1)}{(1-\frac{\xi_{r}\chi_{r}\rho^{*}}{2v_{0}})^{2}}}\Big{]}$ (20)
Let us consider the negative solution $\chi_{t}<\chi_{t}^{*}$. Let us further study the limit $\frac{\chi_{r}}{v_{0}}<<1$. Using the definitions for $l_{r}$ and $l_{t}$, the desired inequality is thus
@@
We can identify this to have the same functional form as Eq.(8) defining the transition line. Equating with Eq. (8), we have that
@@
We can thus infer the values of $\xi_{r}\approx 2c_{1}\approx 0.026$ and $\xi_{0}\approx 1.6\times 10^{-3}$. We thus conclude that the noiseless limit of a minimal hydrodynamic model captures the destabilization transition of the flock. In addition, such transitions within between order/disorder phases (e.g. Fig. 3(b)) are also captured, although there is no explicit spatial-structure instability encoded in the hydrodynamic model (unlike for instance in [37]).
V Conclusions